OpenStudy (anonymous):

If you were asked to convert 25 mg to the unit hg, which of the following would be the first fraction used in the conversion? ten to the negative third power hg over one g ten to the negative third power mg over one hg ten to the negative third power mg over one g ten to the negative third power g over one mg

OpenStudy (anonymous):

@kawaiicat123 @lolly @linn99123 @live4him112 hellpppp

OpenStudy (anonymous):

@ShayleeCutie help

OpenStudy (anonymous):

IM HERE

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

@gold_grace123 help

OpenStudy (anonymous):

helppp @Ail.A @sunnycali @KissesFromKristine @Elsa213 help me please

Elsa213 (elsa213):

your first step is to get rid of the milli (m) prefix. An easy way to do this is to remember that m = 10^-3. Therefore 25 mg = 25 x 10^-3 g = 2.5 x 10^-2 g. Now you need to get to hectograms (why? that's a weird unit!) Since h = 10^2, you need to rewrite your measurement so there is a 10^2 in front of the g: 2.5 x 10^-2 g = 2.5 x 10^-2 x (10^-2 x 10^2) g = 2.5 x 10^-4 x 10^2 g. Now just replace the 10^2 g with hg, and you get 2.5 x 10^-4 hg. That way works even in your head, but the more formal way uses units ratios, in which you continually multiply by 1 (which doesn't change the measurement), using various ways of writing 1. For example, you can write 1 as 10/10 or 123.4/123.4 or (which is relevant here) 10^3 mg / 1 g or 10^2 g / 1 hg. To get rid of the mg, multiply by 1 in such a way that the mg cancel: 25 mg x (1 g / 10^3 mg) = 25 x 10^-3 mg = 2.5 x 10^-2 g. That's the answer you want. To continue on to get the hg, multiply by 1 in such a way that the g cancel: 2.5 x 10^-2 g x (1 hg / 10^2 g) = 2.5 x 10^-4 hg.

OpenStudy (anonymous):

I don't understand